OpenStudy (anonymous):

2 over 2x equals x plus 1

OpenStudy (anonymous):

\[ \frac2{2x}=x+1\\ \frac1x=x+1\\ 1=x^2+x\\ x^2+x-1=0\\ x=\frac{-1\pm\sqrt{5}}{2} \]